### v1

# Symbolic Learning of Interpretable Reduced-Order Models for Jumping Quadruped Robots

Abstract
Reduced-order models are essential for motion planning and control of quadruped robots, as they simplify complex dynamics while preserving critical behaviors. This paper introduces a novel methodology for deriving such interpretable dynamic models, specifically for jumping. We capture the high-dimensional, nonlinear jumping dynamics in a low-dimensional latent space by proposing a learning architecture combining Sparse Identification of Nonlinear Dynamics (SINDy) with physical structural priors on the jump dynamics. Our approach demonstrates superior accuracy to the traditional actuated Spring-loaded Inverted Pendulum (aSLIP) model and is validated through simulation and hardware experiments across different jumping strategies.
keywords:
Quadruped Robots, Dynamic Modeling, Symbolic Regression, Autoencoders, Interpretability, Reduced-order Models.
thanks: This work is supported by the European Union’s Horizon Europe Program Project EMERGE - Grant Agreement No. 101070918. thanks: These authors contributed equally to this work.
1 Introduction
Recently, quadruped robots have garnered significant attention due to their versatility and potential applications in various fields, including search and rescue missions, agricultural tasks, and exploration in hazardous environments (Gong et al., 2023; Sombolestan and Nguyen, 2024). The performance of these robots depends on accurate dynamic modeling to capture their complex locomotion behaviors.
Traditional modeling techniques, such as the Lagrangian and Newton-Euler methods, have been widely used to derive high-fidelity whole-body dynamic models for quadruped robots (Bellicoso et al., 2017; Di Carlo et al., 2018; Yan et al., 2021; Ding et al., 2024b). However, these high-dimensional models are often computationally intractable for real-time applications due to their complexity (Li et al., 2023). To address this, Researchers have developed reduced-order template models that approximate essential dynamics while significantly lowering computational costs. Examples include the famous Spring-Loaded Inverted Pendulum (SLIP) model (Blickhan, 1989) and the Single Rigid Body (SRB) (Hong et al., 2020), which simplifies multi-body interactions into low-dimensional representations, enabling efficient analysis of gaits like running and hopping. However, such simplifications often sacrifice accuracy, particularly in scenarios involving multi-contact transitions, nonlinear limb dynamics, or complex terrain interactions (Spröwitz et al., 2013; Carpentier and Mansard, 2018).
To tackle this issue, we investigate the possibility of using a data-driven strategy to learn reduced-order models directly, with a particular interest in the compact, interpretable, and precise description of jumping dynamics, which is very challenging. Numerous studies have utilized machine learning and deep learning methods to model the complex dynamics of various systems (Calandra et al., 2015; Ogunmolu et al., 2016; Rajendra and Brahmajirao, 2020; Hashemi et al., 2023). Dimensionality reduction techniques like autoencodershave been employed to extract low-dimensional features in a latent space (Hinton and Salakhutdinov, 2006; Gonzalez and Balajewicz, 2018; Refinetti and Goldt, 2022), and then Multilayer Perceptrons (MLPs) or Recurrent Neural Networks (RNNs) can learn the dynamics in this low-dimensional space, but at the expense of interpretability in exchange for accuracy. Recent hybrid approaches that combine physics with learning (Stölzle, 2025), such as Physics-Informed Neural Networks (PINNs) (Raissi et al., 2019), Deep Lagrangian Networks (DeLaNs) (Lutter et al., 2019; Liu et al., 2024), Coupled Oscillator Networks (CONs) (Stölzle and Santina, 2024), integrate physical laws with deep learning, enhancing accuracy, interpretability, data efficiency, and out-of-distribution performance. However, PINNs often require physics knowledge that is only available to domain experts. Furthermore, nonlinear time-dependent behaviors, as for example experienced during contact phases with the environment, are usually not included in the physical priors and can, therefore, not be learned by the DeLaNs. Symbolic regression, particularly SINDy (Brunton et al., 2016; Pandey et al., 2024), presents an alternative through interpretable symbolic equations, though their application to quadruped robots remains unexplored.
This work proposes a novel methodology to derive reduced-order, interpretable dynamic models for jumping robots. Building upon SINDy autoencoders (Champion et al., 2019), our approach uses a linear  [^1] The linearity of the autoencoder is critical for preserving interpretability and facilitating the closed-form, analytical (i.e., without autodifferentiation) computation of time derivatives. autoencoder with shared weights to reduce the dimension of the state space while keeping configurations and maintaining their derivatives distinct in the latent space. We map inputs to latent space through the trained autoencoder, optimizing its weights concurrently with the SINDy latent model using an ad-hoc loss function. Considering the hybrid nature of the system during different jumping phases, we learn a consistent model across various hybrid phases by first training the encoder in one phase (e.g., the launching and landing phase), then optimizing the decoder across all phases, and finally learning the SINDy parameters with frozen autoencoder weights.
The main contributions of this work are threefold:
  • We present a novel modeling approach that integrates SINDy with a linear autoencoder, utilizing shared weights to derive reduced-order, interpretable models for quadruped robots. This method enhances computational efficiency and model interpretability.
  • We propose a sequential training strategy to accurately model the distinct jump phases, ensuring a consistent model across various hybrid phases.
  • We evaluate our novel approach on various quadruped robots with different jumping modalities to demonstrate versatility.
The rest of this paper follows: Section 2 covers preliminary concepts; Section 3.1 discusses the proposed learning architecture; Section 3 describes the remaining aspects of our experimental methodology; Section 4 evaluates the method, featuring simulation results for various quadruped models, latent dimensionality analysis, and hardware validation. Section 5 concludes this work.
2 Preliminary
Traditional modeling approaches, such as multi-rigid-body dynamics, represent quadruped robots as a movable base (i.e., the body) with kinematic chains (i.e., the legs) attached, applicable to various quadruped robots, as shown in Fig. 1.
Figure 1: Quadruped robot with labeled joints and body frame.
As illustrated in Fig. 1, the Unitree Go1, for example, exhibits 121212 Degrees of Freedom (DOFs) for the limbs alone, with 666 underactuated DoFs for the floating base’s position and orientation relative to an inertial frame (Righetti et al., 2011).
The configuration space of quadruped robots is typically parameterized as (Mistry et al., 2010)
$\mathbf{q}=\begin{bmatrix}\mathbf{q}_{\mathrm{j}}^{\top}&\mathbf{q}_{\mathrm{b}}^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{m+6},$ (1)
where $\mathbf{q}_{\mathrm{j}}\in\mathbb{R}^{m}$ contains the coordinates of the actuated joints, with $m$ representing the number of joints, and $\mathbf{q}_{\mathrm{b}}\in SE(3)$ captures the coordinates of the unactuated floating base, including Cartesian positions $(x,y,z)\in\mathbb{R}^{3}$ and Euler angles $(\phi,\theta,\psi)\in SO(3)$. The constrained dynamics follow
$\mathbf{M}(\mathbf{q})\,\ddot{\mathbf{q}}+\mathbf{h}(\mathbf{q},\dot{\mathbf{q}})=\mathbf{S}^{\top}\bm{\tau}+\mathbf{J}_{\tt C}^{\top}(\mathbf{q})\,\bm{\lambda},$ (2)
where $\mathbf{M}(\mathbf{q})\in\mathbb{R}^{(m+6)\times(m+6)}$ is the mass matrix, $\mathbf{h}(\mathbf{q},\dot{\mathbf{q}})\in\mathbb{R}^{m+6}$ captures centripetal, Coriolis, and gravitational forces, $\mathbf{S}=\begin{bmatrix}\bm{I}_{m\times m}&\bm{0}_{m\times 6}\end{bmatrix}$ is the selection matrix for actuated joints, $\bm{\tau}\in\mathbb{R}^{m}$ denotes actuation torques, $\mathbf{J}_{\tt C}\in\mathbb{R}^{k\times(m+6)}$ is the Jacobian matrix for $k$ linearly independent constraints, and $\bm{\lambda}\in\mathbb{R}^{k}$ represents the constraint forces.
Consider a quadruped robot where ground reaction forces (GRFs) constitute the dominant constraint forces acting on the body. For each leg (modeled as an independent kinematic chain), the GRF $\mathbf{F}_{k}\in\mathbb{R}^{3}$ at the $k$-th foot can be computed from the joint torques $\bm{\tau}_{k}$ through the leg Jacobian transpose
$\mathbf{F}_{k}=\mathbf{J}_{k}^{-\top}\bm{\tau}_{k},$ (3)
where $\mathbf{J}_{k}\in\mathbb{R}^{3\times n_{j}}$ is the Jacobian of the $k$-th leg ($n_{j}=m/4$ being the number of joints per leg).
The net effect of GRFs on the robot’s center of mass (CoM) dynamics can be described by the time derivatives of the linear and angular momentum of the CoM
$\dot{\bm{\mathcal{P}}}=\sum_{k=1}^{4}\mathbf{F}_{k},\quad\dot{\bm{L}}=\sum_{k=1}^{4}((\mathbf{p}_{k}-\mathbf{b})\times\mathbf{F}_{k}),$ (4)
where $\bm{\mathcal{P}}\in\mathbb{R}^{3}$ and $\bm{L}\in\mathbb{R}^{3}$ are the linear and angular momentum, respectively. Here, $\mathbf{p}_{k}\in\mathbb{R}^{3}$ denotes the world-frame position of the $k$-th leg, and $\mathbf{b}\in\mathbb{R}^{3}$ is the CoM position in world coordinates.
[…]
which leads to the complete system input formulation
[…]
3 Methodology
3.1 Proposed Learning Architecture
We present a novel framework combining linear autoencoders with SINDy to model quadruped dynamics, incorporating physical priors for dimensionality reduction while preserving essential dynamic features. Our approach achieves dimensionality reduction while preserving the essential dynamic characteristics, enabling efficient representation of complex jumping motions
Fig. 2 provides an overview of the encoded latent space and the established dynamics loop.
Figure 2: Overview of our proposed methodology: We learn symbolic reduced-order dynamics $\ddot{\bm{\xi}}=\bm{\Theta}(\bm{\xi},\dot{\bm{\xi}},\bm{\nu})\,\bm{\Xi}$ that operate in a low-dimensional space $(\bm{\xi},\dot{\bm{\xi}})$. This allows us to predict the time evolution of the higher-dimensional system states $(\mathbf{q},\dot{\mathbf{q}})$. Here, $\mathbf{q}_{0}$ and $\dot{\mathbf{q}}_{0}$ represent the initial state.
Fig. 3 details the training process, including the encoder-decoder structure, latent space dynamics prediction, and real-space dynamics reconstruction. The algorithm includes encoding the initial state into a latent space, predicting dynamics in this latent space, and decoding them back to the full-order system space. We ensure accurate reconstruction of system states and dynamical modeling through specialized loss components.
(a) Training of encoder and decoder in a static setting
(b) Training of symbolic latent dynamics in a dynamic setting
Figure 3: Detailed visualization of the training methodology. (a): The linear autoencoder is first trained on the contact phase dataset, minimizing the reconstruction loss, $\mathcal{L}_{\mathrm{rec}}$, depicted by the black line. Subsequently, in the second step (visualized with a blue line in the flowchart), the encoder weights are kept fixed, and the decoder is fine-tuned on a dataset that includes all motion phases. (b): In the third step, we train, for each motion phase separately, a symbolic dynamical model $\bm{\ddot{\xi}}_{\text{pred}}$ in latent space. Here, we employ the SINDy approach to regress a sparse matrix $\bm{\Xi}$ that selects and scales basis functions from the library $\bm{\Theta}(\bm{\xi},\dot{\bm{\xi}},\bm{\nu})$. The loss components $\mathcal{L}_{\text{SINDy}}$ and $\mathcal{L}_{\text{reg}}$ ensure accurate dynamics prediction and sparsity of the coefficients, respectively. The SINDy loss $\mathcal{L}_{\text{SINDy}}$ is calculated based on the predicted reduced-order model accelerations, $\ddot{\bm{\xi}}_{\mathrm{pred}}$, and the corresponding decoded configuration-space accelerations, $\ddot{\mathbf{q}}_{\mathrm{pred}}$. Importantly, we keep the autoencoder weights frozen, which preserves training stability and ensures that the reconstructions remain accurate during all motion phases.
The linear autoencoder is designed to reduce the dimensionality of the state space while preserving a clear linear relationship with the original space, thereby maintaining direct interpretability. The autoencoder consists of an encoder $\left(\mathbf{W}_{\mathrm{e}},\mathbf{B}_{\mathrm{e}}\right)$ and a decoder $\left(\mathbf{W}_{\mathrm{d}},\mathbf{B}_{\mathrm{d}}\right)$. The encoder’s weights and bias are $\mathbf{W}_{\mathrm{e}}\in\mathbb{R}^{l\times(m+6)}$ and $\mathbf{B}_{\mathrm{e}}\in\mathbb{R}^{l}$, while the decoder’s weights and bias are $\mathbf{W}_{\mathrm{d}}\in\mathbb{R}^{(m+6)\times l}$ and $\mathbf{B}_{\mathrm{d}}\in\mathbb{R}^{m+6}$, where $l$ is the dimension of the desired latent space configuration.
The encoder maps the robot’s configuration $\mathbf{q}$ and its derivatives $\mathbf{q}^{(n)}$ into a lower-dimensional latent space using the linear transformation
$\bm{\xi}^{(n)}=\mathbf{W}_{\mathrm{e}}\,\mathbf{q}^{(n)}+\delta_{n0}\,\mathbf{B}_{\mathrm{e}},\quad n=0,1,2,$ (7)
where $\bm{\xi}\in\mathbb{R}^{l}$ represents the latent representation, $\bm{\xi}^{(n)}$ denotes the $n$-th derivative of $\bm{\xi}$, and $\delta_{n0}$ is the Kronecker delta, equal to 111 when $n=0$ and 0 otherwise. The system’s actuation is mapped into latent space as
$\begin{array}[]{rl}\bm{\nu}&=\mathbf{W}_{\mathrm{e}}^{-\top}\mathbf{u}\in\mathbb{R}^{l}.\end{array}$ (8)
After establishing the latent space mapping, we employ the SINDy algorithm (Brunton et al., 2016) to identify symbolic representations of the latent dynamics (see Fig. 3(b)). SINDy constructs a symbolic library of candidate functions $\Theta(\bm{\xi},\dot{\bm{\xi}},\bm{\nu})$, including polynomial and trigonometric functions of latent variables ($\bm{\xi}$), their derivatives ($\bm{\xi}^{(}n)$), and the transformed control inputs ($\bm{\nu}$). This library represents a comprehensive set of terms that might appear in the system’s governing equations. The objective is to represent the dynamics of the latent variables as a sparse combination of these candidate functions. Mathematically, this is expressed as a sparse regression problem, with the latent dynamics expressed as
$\ddot{\bm{\xi}}_{\mathrm{pred}}=\Theta(\bm{\xi},\dot{\bm{\xi}},\bm{\nu})\,\bm{\Xi},$ (9)
where $\bm{\Xi}\in\mathbb{R}^{p\times l}$ is a sparse matrix of coefficients to be determined, with $p$ being the number of the candidate functions in $\Theta(\bm{\xi},\dot{\bm{\xi}},\bm{\nu})$. The optimization problem is typically solved using sequential thresholding methods to ensure a sparse solution (Brunton et al., 2016; Champion et al., 2019). Promotes sparsity by penalizing the $\ell_{1}$-norm of $\bm{\Xi}$, which encourages most entries in $\Xi$ to be zero. This sparsity-promoting optimization ensures that only a minimal set of candidate functions is selected, resulting in an interpretable and computationally efficient model.
Subsequently, the predicted accelerations in the latent space, $\ddot{\bm{\xi}}_{\mathrm{pred}}$, are transformed back to the original state space by the linear decoder
$\ddot{\mathbf{q}}_{\mathrm{pred}}=\mathbf{W}_{\mathrm{d}}\,\ddot{\bm{\xi}}_{\mathrm{pred}}.$ (10)
For training encoder, decoder, and reduced-order symbolic dynamics, we consider the following loss components:
  • $\mathcal{L}_{\mathrm{recon}}=\left\|\mathbf{q}-\mathbf{q}_{\mathrm{recon}}\right\|_{2}^{2}$ ensures accurate reconstruction of the original state. The reconstructed configuration $\mathbf{q}_{\mathrm{recon}}$ is given by
    $\mathbf{q}_{\mathrm{recon}}=\mathbf{W}_{\mathrm{d}}\,(\mathbf{W}_{\mathrm{e}}\mathbf{q}+\mathbf{B}_{\mathrm{e}})+\mathbf{B}_{\mathrm{d}}.$ (11)
  • $\mathcal{L}_{\mathrm{SINDy}}=\|\ddot{\bm{\xi}}-\ddot{\bm{\xi}}_{\mathrm{pred}}\|_{2}^{2}+\left\|\ddot{\mathbf{q}}-\ddot{\mathbf{q}}_{\mathrm{pred}}\right\|_{2}^{2}$ captures the prediction loss, ensuring that the learned dynamics in the latent space and the reconstructed space are accurate.
  • $\mathcal{L}_{\mathrm{reg}}=\|\bm{\Xi}\|_{1}$ promotes sparsity in SINDy coefficients.
3.2 Training with Multi-Phase Motions
Jumping motion poses unique modeling challenges due to its hybrid dynamical nature. This study considers two types of jumping: the synchronized and the froggy jumps.
In the synchronized jump, the robot launches and lands with full limb contact simultaneously. This motion consists of two dominant phases: (1) a contact phase, where all legs remain grounded during takeoff and landing, and (2) a flight phase, where the robot is entirely airborne.
Conversely, the froggy jump consists of an initial phase where all legs are grounded, followed by a transition phase where only the rear legs remain in contact with the ground, and then progressing to a flight phase.
To effectively model the hybrid jumping dynamics, we propose a structured training methodology (Fig. 3) that combines transfer learning principles (Torrey and Shavlik, 2010) with phase-specific optimization. First, we train the encoder and decoder statically on reconstructing the configurations of the full contact phase by minimizing the $\mathcal{L}_{\mathrm{recon}}$ loss. Then, the encoder weights are frozen, and the decoder is trained on all motion phases, including the flight phase, still minimizing $\mathcal{L}_{\mathrm{recon}}$. Finally, with the encoder and decoder weights fixed, we train the symbolic dynamics for each motion phase (contact, partial contact, and flight) by minimizing the $\mathcal{L}_{\mathrm{SINDy}}$ and $\mathcal{L}_{\mathrm{reg}}$ loss terms, efficiently capturing the hybrid nature of the jumping dynamics.
This training methodology ensures that the final model is physically coherent and can accurately reconstruct the robot’s motion across different phases. By leveraging transfer learning, we create robust models that effectively capture hybrid dynamics, enhance interpretability, and seamlessly handle complex, phase-dependent behaviors.
3.3 Data Collection and Preprocessing
Datasets were collected from both simulated and real-world environment. Simulated data were generated using the PyBullet engine (Coumans and Bai, 2016) with Unitree Go1 and A1 robots. Real-world data were gathered from an actual Go1 robot to validate the simulation results, ensure practical applicability, and benchmark the algorithm’s performance on real-world, noisy data.
Simulation data for the Go1 quadruped robot captured forward jumping motions, as shown in Fig. 4(a), achieved with Model Predictive Control (MPC) (Ding et al., 2024a). Additionally, data from the Unitree A1 quadruped for froggy jumping behavior, as visualized in Fig. 4(b), were collected using a Cartesian PD controller (Nguyen and Nguyen, 2022). For both robots, the dataset includes joint angles and velocities ($\mathbf{q}_{\mathrm{j}},\dot{\mathbf{q}}_{\mathrm{j}}\in\mathbb{R}^{12}$), body position and velocity ($\mathbf{q}_{\mathrm{b}},\dot{\mathbf{q}}_{\mathrm{b}}\in\mathbb{R}^{6}$), input torques ($\bm{\tau}\in\mathbb{R}^{12}$), and the contact state ($\mathbf{c}\in\mathbb{R}^{4}$) for each timestep.
The Go1 simulated data consists of 156 jumps (120 for training, 30 for validation, and 6 for testing), with each jump having 800 data points: 300 for launching, 300 for landing (totaling 600 contact points), and 200 for flight. The A1 simulated dataset contains 150 jumps (100 for training, 40 for validation, and 10 for testing), with each jump comprising 1200 data points: 400 for launch, 400 for landing (800 contact points), and 400 for flight. To enhance robustness, a variant of the datasets with added noise as described in the work from Ding et al. (2024a).
Figure 4: A sequential images of quadruped jumps: (a) A simulated Go1 robot performs a vertically-oriented jump. (b) A simulated A1 executes a froggy jump. (c) A Go1 robot performs a (vertically-oriented) jump in the real world.
Real-world data were collected from a Go1 quadruped, as depicted in Fig. 4(c). The quadruped was operated using the same control algorithm as in the simulation, with data collection performed similarly using the official Go1 SDK. This dataset comprises 15 jumps, distributed as 12 for training, 2 for validation, and 1 for testing, maintaining the same data point distribution as the Go1 simulated data.
[…]
3.4 Evaluation Procedure
[…]
In this study, we regard the aSLIP model (see Appendix.A) as the baseline for predicting the dynamical behavior.
4 Results
This section evaluates the approach using simulated Unitree Go1 and A1 quadrupeds data and real-world data from a Unitree Go1.
4.1 Go1 Simulation Results
4.1.1 Learning in 2D Latent Space for Go1
We initially explored a two-dimensional latent model for the synchronized jumping. Fig. 5 visualizes the transposed weight matrix $\mathbf{W}_{\mathrm{e}}^{\mathrm{T}}$ of the trained encoder, indicating that the (learned) latent dimensions primarily correspond to the robot’s body center positions in the $x$ and $z$ coordinates.
Figure 5: Heatmap of the encoder’s transposed weight matrix $\mathbf{W}_{\mathrm{e}}^{\mathrm{T}}$ trained on the simulated Go1 dataset with two latent dimensions. Brighter colors indicate higher absolute weights. The vertical axis shows the 18 configuration dimensions, and the horizontal axis shows the two latent dimensions.
The learned equations of motion in the latent space for the contact and flight phases, respectively, are
$\begin{cases}\ddot{\xi}_{1}\!\!=\!\!0.36\!-\!0.16\,\dot{\xi}_{1}\!-\!\!0.91\,\dot{\xi}_{2}\!-\!0.75\sin(\dot{\xi}_{1})\!+\!0.11\,\nu_{1}\!-\!0.05\,\nu_{2}\\ \ddot{\xi}_{2}\!\!=\!\!-0.16\!-\!0.23\,\dot{\xi}_{1}\!-\!0.75\,\dot{\xi}_{2}\!-\!0.96\,\sin(\dot{\xi}_{2})\!+\!0.14\,\nu_{2}\end{cases},$ (12)
and
$\begin{cases}\ddot{\xi}_{1}=0.29-1.22\,\dot{\xi}_{1}-1.00\sin(\dot{\xi}_{1})+51.05\,\nu_{1}\\ \ddot{\xi}_{2}=0.42\,\dot{\xi}_{1}+0.52\sin(\dot{\xi}_{1})\end{cases}.$ (13)
The prediction results for simulated Go 1 jumping are shown in Fig. 6.
Figure 6: Evaluation of the trained model with two latent dimensions on the simulated Go1 jumping dataset.
Fig. 7 compares the body’s linear position along the $x$ and $z$ axes, demonstrating that the learned model outperforms, especially in the $x$ axis, the baseline aSLIP model.
Figure 7: Comparison between the learned model and the aSLIP model, on the task of predicting the CoM linear positions along the $x$ and $z$ axes, where the simulated Go 1 performs a jumping motion.
4.1.2 Latent Space Dimensionality Tradeoff Analysis
We analyze the effect of latent dimensionality on model design. As shown in Fig. 8, increasing to 18 dimensions does not change the dominance of key full-order components.
Figure 8: Heatmap visualization of the transposed weight matrix $\mathbf{W}_{\mathrm{e}}^{\mathrm{T}}$ of the encoder trained on the simulated Go1 dataset in the case of 181818 latent dimensions.
Fig.9 further evaluates the trade-off between accuracy (via final reconstruction error $\mathcal{E}_{\text{dec}}$) and complexity (via active SINDy coefficients $|\bm{\Xi}|$) across 18 contact phase models with varying latent dimensions.
Figure 9: The trade-off between accuracy and complexity with increasing latent dimensions evaluated on the simulated Go1 dataset. The error bars depict the standard deviation from five different random seeds.
To quantitatively evaluate this trade-off, we employ a combined loss $\mathcal{L}_{\text{mod}}$ that is inspired by the Akaike Information Criterion (AIC) for model selection (Bozdogan, 1987) and calculated as
$\mathcal{L}_{\text{mod}}=2\,\mathcal{E}_{\text{dec}}+2\,\lambda\,\log{(|\bm{\Xi}|)},$ (14)
with $\lambda=0.001$, balancing the loss terms, the combined loss (red line in Fig.9) suggests optimal performance with 4–6 latent dimensions. This aligns with the visibly significant components in Fig.8, where further increasing dimensionality yields marginal gains at the cost of complexity.
4.2 A1 Simulation Results
We also conducted experiments on data collected with a simulated Unitree A1 quadruped executing a froggy jump to evaluate the algorithm’s effectiveness across different platforms. The results are illustrated in Fig. 10.
Figure 10: Evaluation of the proposed method with four latent dimensions on A1 froggy jumping.
A notable difference is observed in the body angular position $y$ between the Unitree A1 and the Unitree Go1. The froggy jump causes a more pronounced angular rotation along the $y$-axis in the A1’s case. Despite the increased significance of this angular dimension for the jump motion, the four-dimensional model successfully captures this dynamic and most joint positions.
4.3 Go1 Real World Results
Given the scarcity of real-world compared to simulated data, we first train the model on simulation data with added noise and then subsequently fine-tune it on the few-sample real-world dataset. Fig. 11 illustrates the model’s ability to replicate a jump on the hardware, comparing joint angles and CoM positions and orientations.
Figure 11: Evaluation of the learned reduced-order model with four latent dimensions on the Go1 real-world testing dataset.
The proposed model performs well on actual hardware data, despite reduced accuracy during flight and initial landing phases compared to simulations. Nonetheless, the model effectively manages real-world noise and complexity, demonstrating robust performance.
5 Discussion and Conclusion
The results of our symbolic learning framework show that a reduced latent space can effectively preserve key dynamic patterns of the quadruped robot jumping motion. By combining autoencoder-based embeddings with sparse regression, we obtain interpretable and compact models that approximate the complex full-order dynamics. Our analysis of latent dimensionality shows that increasing beyond 4–6 dimensions offers diminishing returns in accuracy while introducing additional model complexity. The sim-to-real transfer results indicate that the model retains core dynamic features and generalizes well despite limited real-world data, validating its robustness to measurement noise and domain shift.
In future work, we aim to extend this methodology to more dynamic behaviors, such as walking and running, and to systems involving rapid contact changes or higher-dimensional morphologies, including humanoids.
Appendix A Baseline: actuated SLIP Model
The actuated SLIP (aSLIP) (Ding et al., 2024a) model extends the basic SLIP model by incorporating actuators, allows for enhanced control over legged jumping, as depicted in Fig. 12.
Figure 12: The aSLIP model. The actuation (denoted as $m\mathbf{u}$) is explicitly captured to enhance controllability.
The CoM acceleration $\mathbf{\ddot{b}}$ is
$\begin{array}[]{r l}\text{In flight: }&\mathbf{\ddot{b}}=\bm{g}\\ \text{In Contact: }&\mathbf{\ddot{b}}=\frac{k_{\mathrm{s}}|\bm{l}_{0}-\bm{l}|\hat{\bm{l}}}{m}+\bm{g}+\bm{u}\end{array}$ (15)
where $k_{\mathrm{s}}\in\mathbb{R}^{+}$ is the spring constant, $m\in\mathbb{R}^{+}$ is the body mass, $\bm{l}=\mathbf{b}-\mathbf{p}_{k}$ is the leg vector, $\hat{\bm{l}}$ is the unit vector along the leg retraction direction, $\bm{l}_{0}$ is the rest length, and $\bm{g}=[0,0,-g]^{\top}$ with $g$ being the vertical gravitational constant. $\bm{u}\in\mathbb{R}^{3}$ are the driving forces generated by joint motors, which can be seen as the accumulated GRFs on four legs, as calculated using (3).

### v2

@@
 Abstract
-Reduced-order models are essential for motion planning and control of quadruped robots, as they simplify complex dynamics while preserving critical behaviors.
-This paper introduces a novel methodology for deriving such interpretable dynamic models, specifically for jumping.
-We capture the high-dimensional, nonlinear jumping dynamics in a low-dimensional latent space by proposing a learning architecture combining Sparse Identification of Nonlinear Dynamics (SINDy) with physical structural priors on the jump dynamics.
-Our approach demonstrates superior accuracy to the traditional actuated Spring-loaded Inverted Pendulum (aSLIP) model and is validated through simulation and hardware experiments across different jumping strategies.
+Reduced-order models are central to motion planning and control of quadruped robots, yet existing templates are often hand-crafted for a specific locomotion modality.
+This motivates the need for automatic methods that extract task-specific, interpretable low-dimensional dynamics directly from data.
+We propose a methodology that combines a linear autoencoder with symbolic regression to derive such models.
+The linear autoencoder provides a consistent latent embedding for configurations, velocities, accelerations, and inputs, enabling the sparse identification of nonlinear dynamics (SINDy) to operate in a compact, physics-aligned space.
+A multi-phase, hybrid-aware training scheme ensures coherent latent coordinates across contact transitions.
+We focus our validation on quadruped jumping—a representative, challenging, yet contained scenario in which a principled template model is especially valuable.
+The resulting symbolic dynamics outperform the state-of-the-art handcrafted actuated spring-loaded inverted pendulum (aSLIP) baseline in simulation and hardware across multiple robots and jumping modalities.
 keywords:
-Quadruped Robots, Dynamic Modeling, Symbolic Regression, Autoencoders, Interpretability, Reduced-order Models.
+Quadruped jumping, Symbolic Regression, Autoencoders, Interpretability, Reduced-order Model.
 thanks: This work is supported by the European Union’s Horizon Europe Program Project EMERGE - Grant Agreement No. 101070918. thanks: These authors contributed equally to this work.
 1 Introduction
-Recently, quadruped robots have garnered significant attention due to their versatility and potential applications in various fields, including search and rescue missions, agricultural tasks, and exploration in hazardous environments (Gong et al., 2023; Sombolestan and Nguyen, 2024).
-The performance of these robots depends on accurate dynamic modeling to capture their complex locomotion behaviors.
-Traditional modeling techniques, such as the Lagrangian and Newton-Euler methods, have been widely used to derive high-fidelity whole-body dynamic models for quadruped robots (Bellicoso et al., 2017; Di Carlo et al., 2018; Yan et al., 2021; Ding et al., 2024b).
-However, these high-dimensional models are often computationally intractable for real-time applications due to their complexity (Li et al., 2023).
-To address this, Researchers have developed reduced-order template models that approximate essential dynamics while significantly lowering computational costs.
-Examples include the famous Spring-Loaded Inverted Pendulum (SLIP) model (Blickhan, 1989) and the Single Rigid Body (SRB) (Hong et al., 2020), which simplifies multi-body interactions into low-dimensional representations, enabling efficient analysis of gaits like running and hopping.
-However, such simplifications often sacrifice accuracy, particularly in scenarios involving multi-contact transitions, nonlinear limb dynamics, or complex terrain interactions (Spröwitz et al., 2013; Carpentier and Mansard, 2018).
-To tackle this issue, we investigate the possibility of using a data-driven strategy to learn reduced-order models directly, with a particular interest in the compact, interpretable, and precise description of jumping dynamics, which is very challenging.
-Numerous studies have utilized machine learning and deep learning methods to model the complex dynamics of various systems (Calandra et al., 2015; Ogunmolu et al., 2016; Rajendra and Brahmajirao, 2020; Hashemi et al., 2023).
-Dimensionality reduction techniques like autoencodershave been employed to extract low-dimensional features in a latent space (Hinton and Salakhutdinov, 2006; Gonzalez and Balajewicz, 2018; Refinetti and Goldt, 2022), and then Multilayer Perceptrons (MLPs) or Recurrent Neural Networks (RNNs) can learn the dynamics in this low-dimensional space, but at the expense of interpretability in exchange for accuracy.
-Recent hybrid approaches that combine physics with learning (Stölzle, 2025), such as Physics-Informed Neural Networks (PINNs) (Raissi et al., 2019), Deep Lagrangian Networks (DeLaNs) (Lutter et al., 2019; Liu et al., 2024), Coupled Oscillator Networks (CONs) (Stölzle and Santina, 2024), integrate physical laws with deep learning, enhancing accuracy, interpretability, data efficiency, and out-of-distribution performance.
-However, PINNs often require physics knowledge that is only available to domain experts.
-Furthermore, nonlinear time-dependent behaviors, as for example experienced during contact phases with the environment, are usually not included in the physical priors and can, therefore, not be learned by the DeLaNs.
-Symbolic regression, particularly SINDy (Brunton et al., 2016; Pandey et al., 2024), presents an alternative through interpretable symbolic equations, though their application to quadruped robots remains unexplored.
-This work proposes a novel methodology to derive reduced-order, interpretable dynamic models for jumping robots.
-Building upon SINDy autoencoders (Champion et al., 2019), our approach uses a linear  [^1] The linearity of the autoencoder is critical for preserving interpretability and facilitating the closed-form, analytical (i.e., without autodifferentiation) computation of time derivatives.
-autoencoder with shared weights to reduce the dimension of the state space while keeping configurations and maintaining their derivatives distinct in the latent space.
-We map inputs to latent space through the trained autoencoder, optimizing its weights concurrently with the SINDy latent model using an ad-hoc loss function.
-Considering the hybrid nature of the system during different jumping phases, we learn a consistent model across various hybrid phases by first training the encoder in one phase (e.g., the launching and landing phase), then optimizing the decoder across all phases, and finally learning the SINDy parameters with frozen autoencoder weights.
-The main contributions of this work are threefold:
-  • We present a novel modeling approach that integrates SINDy with a linear autoencoder, utilizing shared weights to derive reduced-order, interpretable models for quadruped robots.
-This method enhances computational efficiency and model interpretability.
-  • We propose a sequential training strategy to accurately model the distinct jump phases, ensuring a consistent model across various hybrid phases.
-  • We evaluate our novel approach on various quadruped robots with different jumping modalities to demonstrate versatility.
-The rest of this paper follows: Section 2 covers preliminary concepts; Section 3.1 discusses the proposed learning architecture; Section 3 describes the remaining aspects of our experimental methodology; Section 4 evaluates the method, featuring simulation results for various quadruped models, latent dimensionality analysis, and hardware validation.
-Section 5 concludes this work.
+Legged robots are now pervasive systems, deployed in applications ranging from planetary exploration (Arm et al., 2023) to search-and-rescue operations (Bai et al., 2018).
+Yet, despite this growing presence, their dynamics remain remarkably difficult to model—especially when seeking a description expressive enough to capture essential behaviors and lightweight enough for real-time use.
+Classic Lagrangian and Newton–Euler formulations provide full-order models of legged systems (Bellicoso et al., 2017; Di Carlo et al., 2018; Yan et al., 2021; Ding et al., 2024b), but these models remain too computationally heavy for online deployment (Li et al., 2023).
+This has led to reduced-order templates such as spring-loaded inverted pendulum (SLIP) (Blickhan, 1989) and single rigid body dynamics (SRBD) (Hong et al., 2020), which capture dominant behaviors at far lower cost.
+Yet these templates lose accuracy in multi-contact transitions, rapid reconfiguration, and complex terrain (Carpentier and Mansard, 2018).
+In practice, one must choose between high-dimensional models that are difficult to deploy and compact, task-specific heuristics.
+A principled way to automatically extract reduced-order models—accurate, compact, and interpretable—would significantly advance the analysis and control of hybrid locomotion dynamics.
+However, such a method is still missing.
+In this paper, we turn to interpretable data-driven techniques to address this challenge.
+Data-driven learning has shown strong capabilities in modeling physical systems, with numerous works leveraging neural networks to capture complex dynamics (Calandra et al., 2015; Ogunmolu et al., 2016; Rajendra and Brahmajirao, 2020; Hashemi et al., 2023).
+Dimensionality-reduction methods such as autoencoders extract low-dimensional latent features (Hinton and Salakhutdinov, 2006; Gonzalez and Balajewicz, 2018; Refinetti and Goldt, 2022), after which multi-layer perceptrons (MLPs) or recurrent neural networks (RNNs) can model reduced-order dynamics—often gaining accuracy at the expense of interpretability.
+Physics-informed approaches, including physics-informed neural networks (PINNs) (Raissi et al., 2019), deep Lagrangian networks (DeLaN) (Lutter et al., 2019; Liu et al., 2024), and coupled oscillator networks (Stölzle and Santina, 2024), improve interpretability, data efficiency, and out-of-distribution performance.
+Still, none of these methods handle the hybrid, time-dependent structure inherent to locomotion.
+This limitation is particularly acute for floating-base dynamics, where contacts induce abrupt changes in inertial and potential forces—effects that current learning frameworks are not designed to capture.
+Given this gap, no existing work provides a systematic method for learning reduced-order models of legged systems.
+We address this by combining linear autoencoders with symbolic regression, specifically the sparse identification of nonlinear dynamics (SINDy) (Brunton et al., 2016; Pandey et al., 2024), extending prior autoencoder-based formulations (Champion et al., 2019) to quadruped dynamics, where such techniques remain unexplored.
+The linear autoencoder is essential: its constant Jacobian—equal to its weight matrix—provides a unified linear mapping of configurations, velocities, accelerations, and inputs, enabling consistent and interpretable identification of reduced-order dynamics.
+Direct identification in the observation space is impractical due to scale disparities and dimensionality, while nonlinear autoencoders introduce state-dependent latent geometries and distort physical structure through separate nonlinear mappings of positions and velocities.
+To obtain a representation valid across hybrid phases, we adopt a multi-phase training scheme: train the encoder in the contact phase (launch + landing), refine the decoder across all phases, and finally identify SINDy parameters with the autoencoder frozen.
+We validate the approach on jumping motions, a demanding form of dynamic locomotion that enables navigation in complex environments (Nguyen et al., 2019; Atanassov et al., 2024; Ding et al., 2023) yet remains challenging due to high accelerations and impact-driven dynamics.
+In summary, this work contributes to legged robotics and physics-informed machine learning through:
+  • A physics-aligned linear latent embedding that provides consistent linear mappings of positions, velocities, accelerations, and inputs, enabling symbolic learning of reduced-order and interpretable jumping dynamics.
+  • A sequential multi-phase training strategy that preserves coherent latent coordinates and symbolic models across the hybrid phases of a jump.
+  • The first demonstration of learning reduced-order dynamics for legged locomotion, with evaluation across different quadruped robots and jumping modalities.
 2 Preliminary
-Traditional modeling approaches, such as multi-rigid-body dynamics, represent quadruped robots as a movable base (i.e., the body) with kinematic chains (i.e., the legs) attached, applicable to various quadruped robots, as shown in Fig.
-1.
+2.1 Robot description
+Traditional modeling approaches represent the quadruped robot as a movable base (i.e., the body) with kinematic chains (i.e., the legs) attached, as shown in Fig. 1.
 Figure 1: Quadruped robot with labeled joints and body frame.
 As illustrated in Fig.
-1, the Unitree Go1, for example, exhibits 121212 Degrees of Freedom (DOFs) for the limbs alone, with 666 underactuated DoFs for the floating base’s position and orientation relative to an inertial frame (Righetti et al., 2011).
+1, the Unitree Go1, for example, exhibits $12$ degrees of freedom (DOFs) for the limbs alone, with $6$ underactuated DoFs for the floating base’s position and orientation relative to an inertial frame (Righetti et al., 2011).
 The configuration space of quadruped robots is typically parameterized as (Mistry et al., 2010)
@@
 […]
-3 Methodology
-3.1 Proposed Learning Architecture
-We present a novel framework combining linear autoencoders with SINDy to model quadruped dynamics, incorporating physical priors for dimensionality reduction while preserving essential dynamic features.
-Our approach achieves dimensionality reduction while preserving the essential dynamic characteristics, enabling efficient representation of complex jumping motions
-Fig. 2 provides an overview of the encoded latent space and the established dynamics loop.
-Figure 2: Overview of our proposed methodology: We learn symbolic reduced-order dynamics $\ddot{\bm{\xi}}=\bm{\Theta}(\bm{\xi},\dot{\bm{\xi}},\bm{\nu})\,\bm{\Xi}$ that operate in a low-dimensional space $(\bm{\xi},\dot{\bm{\xi}})$.
-This allows us to predict the time evolution of the higher-dimensional system states $(\mathbf{q},\dot{\mathbf{q}})$.
-Here, $\mathbf{q}_{0}$ and $\dot{\mathbf{q}}_{0}$ represent the initial state.
-Fig.
-3 details the training process, including the encoder-decoder structure, latent space dynamics prediction, and real-space dynamics reconstruction.
-The algorithm includes encoding the initial state into a latent space, predicting dynamics in this latent space, and decoding them back to the full-order system space.
-We ensure accurate reconstruction of system states and dynamical modeling through specialized loss components.
-(a) Training of encoder and decoder in a static setting
-(b) Training of symbolic latent dynamics in a dynamic setting
-Figure 3: Detailed visualization of the training methodology.
-(a): The linear autoencoder is first trained on the contact phase dataset, minimizing the reconstruction loss, $\mathcal{L}_{\mathrm{rec}}$, depicted by the black line.
-Subsequently, in the second step (visualized with a blue line in the flowchart), the encoder weights are kept fixed, and the decoder is fine-tuned on a dataset that includes all motion phases.
-(b): In the third step, we train, for each motion phase separately, a symbolic dynamical model $\bm{\ddot{\xi}}_{\text{pred}}$ in latent space.
-Here, we employ the SINDy approach to regress a sparse matrix $\bm{\Xi}$ that selects and scales basis functions from the library $\bm{\Theta}(\bm{\xi},\dot{\bm{\xi}},\bm{\nu})$.
-The loss components $\mathcal{L}_{\text{SINDy}}$ and $\mathcal{L}_{\text{reg}}$ ensure accurate dynamics prediction and sparsity of the coefficients, respectively.
-The SINDy loss $\mathcal{L}_{\text{SINDy}}$ is calculated based on the predicted reduced-order model accelerations, $\ddot{\bm{\xi}}_{\mathrm{pred}}$, and the corresponding decoded configuration-space accelerations, $\ddot{\mathbf{q}}_{\mathrm{pred}}$.
-Importantly, we keep the autoencoder weights frozen, which preserves training stability and ensures that the reconstructions remain accurate during all motion phases.
-The linear autoencoder is designed to reduce the dimensionality of the state space while preserving a clear linear relationship with the original space, thereby maintaining direct interpretability.
-The autoencoder consists of an encoder $\left(\mathbf{W}_{\mathrm{e}},\mathbf{B}_{\mathrm{e}}\right)$ and a decoder $\left(\mathbf{W}_{\mathrm{d}},\mathbf{B}_{\mathrm{d}}\right)$.
-The encoder’s weights and bias are $\mathbf{W}_{\mathrm{e}}\in\mathbb{R}^{l\times(m+6)}$ and $\mathbf{B}_{\mathrm{e}}\in\mathbb{R}^{l}$, while the decoder’s weights and bias are $\mathbf{W}_{\mathrm{d}}\in\mathbb{R}^{(m+6)\times l}$ and $\mathbf{B}_{\mathrm{d}}\in\mathbb{R}^{m+6}$, where $l$ is the dimension of the desired latent space configuration.
-The encoder maps the robot’s configuration $\mathbf{q}$ and its derivatives $\mathbf{q}^{(n)}$ into a lower-dimensional latent space using the linear transformation
-$\bm{\xi}^{(n)}=\mathbf{W}_{\mathrm{e}}\,\mathbf{q}^{(n)}+\delta_{n0}\,\mathbf{B}_{\mathrm{e}},\quad n=0,1,2,$ (7)
-where $\bm{\xi}\in\mathbb{R}^{l}$ represents the latent representation, $\bm{\xi}^{(n)}$ denotes the $n$-th derivative of $\bm{\xi}$, and $\delta_{n0}$ is the Kronecker delta, equal to 111 when $n=0$ and 0 otherwise.
-The system’s actuation is mapped into latent space as
-$\begin{array}[]{rl}\bm{\nu}&=\mathbf{W}_{\mathrm{e}}^{-\top}\mathbf{u}\in\mathbb{R}^{l}.\end{array}$ (8)
-After establishing the latent space mapping, we employ the SINDy algorithm (Brunton et al., 2016) to identify symbolic representations of the latent dynamics (see Fig.
-3(b)).
-SINDy constructs a symbolic library of candidate functions $\Theta(\bm{\xi},\dot{\bm{\xi}},\bm{\nu})$, including polynomial and trigonometric functions of latent variables ($\bm{\xi}$), their derivatives ($\bm{\xi}^{(}n)$), and the transformed control inputs ($\bm{\nu}$).
+2.2 Encoder–Decoder Representation Learning
+The autoencoder, i.e., encoder–decoder architecture, enables one to obtain a compact but informative representation of high-dimensional state variables (Hinton and Salakhutdinov, 2006).
+The encoder $E(\cdot)$ maps the original variable $\bm{x}\in\mathbb{R}^{n}$ into a low-dimensional latent vector $\bm{z}=E(\bm{x})\in\mathbb{R}^{k}$ with $k\ll n$, effectively capturing the essential features that characterize the system’s dynamics.
+The decoder $D(\cdot)$ reconstructs the high-dimensional signal from the latent representation, i.e., $\hat{\bm{x}}=D(\bm{z})$, ensuring that the latent space preserves the underlying structure of the original data.
+This framework not only reduces computational complexity but also provides a smooth and continuous latent manifold that is amenable to downstream tasks such as policy learning.
+2.3 SINDy
+As a typical symbolic learning approach, SINDy (Brunton et al., 2016) enables us to identify symbolic representations of the dynamics by constructing a symbolic library of candidate functions $\Theta(\bm{\xi},\dot{\bm{\xi}},\bm{\nu})$, including polynomial and trigonometric functions of latent variables ($\bm{\xi}$), their derivatives ($\bm{\xi}^{(n)}$), and the transformed control inputs ($\bm{\nu}$).
 This library represents a comprehensive set of terms that might appear in the system’s governing equations.
@@
 Promotes sparsity by penalizing the $\ell_{1}$-norm of $\bm{\Xi}$, which encourages most entries in $\Xi$ to be zero.
-This sparsity-promoting optimization ensures that only a minimal set of candidate functions is selected, resulting in an interpretable and computationally efficient model.
-Subsequently, the predicted accelerations in the latent space, $\ddot{\bm{\xi}}_{\mathrm{pred}}$, are transformed back to the original state space by the linear decoder
+This sparsity-promoting optimization ensures that only a minimal set of candidate functions is selected, resulting in an interpretable model.
+3 Methodology
+3.1 Proposed Learning Architecture
+We combine linear autoencoders with SINDy to model jumping dynamics in the latent space, achieving dimensionality reduction while preserving the essential dynamic characteristics, enabling efficient representation of complex jumping motions.
+Fig.
+2 provides an overview of the encoded latent space and the established dynamics loop.
+Figure 2: Methodology overview: Symbolic latent-space dynamics $\ddot{\bm{\xi}}=\bm{\Theta}(\bm{\xi},\dot{\bm{\xi}},\bm{\nu})\,\bm{\Xi}$ are learned to predict the evolution of the full-state system from $(\mathbf{q},\dot{\mathbf{q}})$.
+Fig.
+3 details the core components, including the encoder-decoder structure, latent space dynamics prediction, and real-space dynamics reconstruction.
+That is, we encode the initial state into a latent space, predict the dynamics in this latent space, and decode them back to the full-order system space.
+In particular, we ensure accurate reconstruction of system states and dynamical modeling through specialized loss components.
+(a) Training of encoder and decoder in a static setting
+(b) Training of symbolic latent dynamics in a dynamic setting
+Figure 3: Flowchart of the pipeline.
+(a) The linear autoencoder is first trained on the contact phase dataset, minimizing the reconstruction loss, $\mathcal{L}_{\mathrm{rec}}$, depicted by the black line.
+Subsequently, by fixing the encoder, the decoder is fine-tuned on a dataset that includes all motion phases (see the the blue line).
+(b) In the third step, we train, for each motion phase separately, a symbolic dynamical model $\bm{\ddot{\xi}}_{\text{pred}}$ in latent space.
+Here, we employ the SINDy approach to regress a sparse matrix $\bm{\Xi}$ that selects and scales basis functions from the library $\bm{\Theta}(\bm{\xi},\dot{\bm{\xi}},\bm{\nu})$.
+The loss components $\mathcal{L}_{\text{SINDy}}$ and $\mathcal{L}_{\text{reg}}$ ensure accurate dynamics prediction and sparsity of the coefficients, respectively.
+The SINDy loss $\mathcal{L}_{\text{SINDy}}$ is calculated based on the predicted reduced-order model accelerations, $\ddot{\bm{\xi}}_{\mathrm{pred}}$, and the corresponding decoded configuration-space accelerations, $\ddot{\mathbf{q}}_{\mathrm{pred}}$.
+To maintain direct interpretability, we adopt the linear autoencoder to reduce the dimensionality of the state space while preserving a linear relationship with the original space.
+The autoencoder consists of an encoder $\left(\mathbf{W}_{\mathrm{e}},\mathbf{B}_{\mathrm{e}}\right)$ and a decoder $\left(\mathbf{W}_{\mathrm{d}},\mathbf{B}_{\mathrm{d}}\right)$.
+The encoder’s weights and bias are $\mathbf{W}_{\mathrm{e}}\in\mathbb{R}^{l\times(m+6)}$ and $\mathbf{B}_{\mathrm{e}}\in\mathbb{R}^{l}$, while the decoder’s weights and bias are $\mathbf{W}_{\mathrm{d}}\in\mathbb{R}^{(m+6)\times l}$ and $\mathbf{B}_{\mathrm{d}}\in\mathbb{R}^{m+6}$, where $l$ is the dimension of the desired latent space.
+With this encoder, we map the robot’s configuration $\mathbf{q}$ and its derivatives $\mathbf{q}^{(n)}$ into a lower-dimensional latent space using the linear transformation
+$\bm{\xi}^{(n)}=\mathbf{W}_{\mathrm{e}}\,\mathbf{q}^{(n)}+\delta_{n0}\,\mathbf{B}_{\mathrm{e}},\quad n=0,1,2,$ (8)
+where $\bm{\xi}\in\mathbb{R}^{l}$ represents the latent representation, $\bm{\xi}^{(n)}$ denotes the $n$-th derivative of $\bm{\xi}$, and $\delta_{n0}$ is the Kronecker delta, equal to $1$ when $n=0$ and 0 otherwise.
+The system’s actuation is mapped into latent space as
+$\begin{array}[]{rl}\bm{\nu}&=\mathbf{W}_{\mathrm{e}}^{-\top}\mathbf{u}\in\mathbb{R}^{l}.\end{array}$ (9)
+After establishing the latent space mapping, we employ the SINDy algorithm to identify symbolic representations of the latent dynamics (see Fig.
+3(b)).
+Particularly, we follow Eq.
+(7) to predict the accelerations.
+Subsequently, the predicted accelerations $\ddot{\bm{\xi}}_{\mathrm{pred}}$ are transformed back to the original state space by the linear decoder
 $\ddot{\mathbf{q}}_{\mathrm{pred}}=\mathbf{W}_{\mathrm{d}}\,\ddot{\bm{\xi}}_{\mathrm{pred}}.$ (10)
-For training encoder, decoder, and reduced-order symbolic dynamics, we consider the following loss components:
+For training the encoder, decoder, and reduced-order symbolic dynamics, we consider the following loss components:
   • $\mathcal{L}_{\mathrm{recon}}=\left\|\mathbf{q}-\mathbf{q}_{\mathrm{recon}}\right\|_{2}^{2}$ ensures accurate reconstruction of the original state.
@@
   • $\mathcal{L}_{\mathrm{reg}}=\|\bm{\Xi}\|_{1}$ promotes sparsity in SINDy coefficients.
-3.2 Training with Multi-Phase Motions
-Jumping motion poses unique modeling challenges due to its hybrid dynamical nature. This study considers two types of jumping: the synchronized and the froggy jumps.
-In the synchronized jump, the robot launches and lands with full limb contact simultaneously.
+3.2 Multi-Phase Training Strategy
+Jumping motion poses unique modeling challenges due to its hybrid dynamical nature. This study considers two types of jumping: the pronking and the froggy jumps.
+In the pronking jump, the robot launches and lands with full limb contact simultaneously.
 This motion consists of two dominant phases: (1) a contact phase, where all legs remain grounded during takeoff and landing, and (2) a flight phase, where the robot is entirely airborne.
 Conversely, the froggy jump consists of an initial phase where all legs are grounded, followed by a transition phase where only the rear legs remain in contact with the ground, and then progressing to a flight phase.
-To effectively model the hybrid jumping dynamics, we propose a structured training methodology (Fig.
+To effectively model the hybrid dynamics, we propose a structured training methodology (Fig.
 3) that combines transfer learning principles (Torrey and Shavlik, 2010) with phase-specific optimization.
 First, we train the encoder and decoder statically on reconstructing the configurations of the full contact phase by minimizing the $\mathcal{L}_{\mathrm{recon}}$ loss.
-Then, the encoder weights are frozen, and the decoder is trained on all motion phases, including the flight phase, still minimizing $\mathcal{L}_{\mathrm{recon}}$.
-Finally, with the encoder and decoder weights fixed, we train the symbolic dynamics for each motion phase (contact, partial contact, and flight) by minimizing the $\mathcal{L}_{\mathrm{SINDy}}$ and $\mathcal{L}_{\mathrm{reg}}$ loss terms, efficiently capturing the hybrid nature of the jumping dynamics.
-This training methodology ensures that the final model is physically coherent and can accurately reconstruct the robot’s motion across different phases.
-By leveraging transfer learning, we create robust models that effectively capture hybrid dynamics, enhance interpretability, and seamlessly handle complex, phase-dependent behaviors.
+Then, the encoder weights are frozen, and the decoder is trained on all motion phases, including the flight phase, while still minimizing $\mathcal{L}_{\mathrm{recon}}$.
+Finally, with the encoder and decoder weights fixed, we train the symbolic dynamics for each motion phase (contact, partial contact, and flight) by minimizing the $\mathcal{L}_{\mathrm{SINDy}}$ and $\mathcal{L}_{\mathrm{reg}}$ loss terms.
+In this way, the final model is physically coherent and can accurately reconstruct the robot’s motion across different phases.
 3.3 Data Collection and Preprocessing
-Datasets were collected from both simulated and real-world environment.
+Datasets were collected from both simulated and real-world environments.
 Simulated data were generated using the PyBullet engine (Coumans and Bai, 2016) with Unitree Go1 and A1 robots.
-Real-world data were gathered from an actual Go1 robot to validate the simulation results, ensure practical applicability, and benchmark the algorithm’s performance on real-world, noisy data.
-Simulation data for the Go1 quadruped robot captured forward jumping motions, as shown in Fig.
-4(a), achieved with Model Predictive Control (MPC) (Ding et al., 2024a).
-Additionally, data from the Unitree A1 quadruped for froggy jumping behavior, as visualized in Fig.
-4(b), were collected using a Cartesian PD controller (Nguyen and Nguyen, 2022).
-For both robots, the dataset includes joint angles and velocities ($\mathbf{q}_{\mathrm{j}},\dot{\mathbf{q}}_{\mathrm{j}}\in\mathbb{R}^{12}$), body position and velocity ($\mathbf{q}_{\mathrm{b}},\dot{\mathbf{q}}_{\mathrm{b}}\in\mathbb{R}^{6}$), input torques ($\bm{\tau}\in\mathbb{R}^{12}$), and the contact state ($\mathbf{c}\in\mathbb{R}^{4}$) for each timestep.
-The Go1 simulated data consists of 156 jumps (120 for training, 30 for validation, and 6 for testing), with each jump having 800 data points: 300 for launching, 300 for landing (totaling 600 contact points), and 200 for flight.
-The A1 simulated dataset contains 150 jumps (100 for training, 40 for validation, and 10 for testing), with each jump comprising 1200 data points: 400 for launch, 400 for landing (800 contact points), and 400 for flight.
-To enhance robustness, a variant of the datasets with added noise as described in the work from Ding et al.
-(2024a).
-Figure 4: A sequential images of quadruped jumps: (a) A simulated Go1 robot performs a vertically-oriented jump.
-(b) A simulated A1 executes a froggy jump.
-(c) A Go1 robot performs a (vertically-oriented) jump in the real world.
-Real-world data were collected from a Go1 quadruped, as depicted in Fig.
+Real-world data were gathered from an actual Go1 robot.
+Simulation data for Go1 captured forward pronking motions (see Fig.
+4(a)), achieved with model predictive control (MPC) (Ding et al., 2024a).
+Data from A1 for froggy jumping (see Fig.
+4(b)) were collected using a Cartesian PD controller (Nguyen and Nguyen, 2022).
+The dataset includes joint angles and velocities ($\mathbf{q}_{\mathrm{j}},\dot{\mathbf{q}}_{\mathrm{j}}\in\mathbb{R}^{12}$), body position and velocity ($\mathbf{q}_{\mathrm{b}},\dot{\mathbf{q}}_{\mathrm{b}}\in\mathbb{R}^{6}$), input torques ($\bm{\tau}\in\mathbb{R}^{12}$), and the contact state ($\mathbf{c}\in\mathbb{R}^{4}$) for each timestep.
+The Go1 simulated data consists of 156 jumps (120 for training, 30 for validation, and 6 for testing), with each jump having 800 points: 300 for launching, 300 for landing, and 200 for flight.
+The A1 dataset contains 150 jumps (100 for training, 40 for validation, and 10 for testing), with each jump comprising 1200 points: 400 for launch, 400 for landing, and 400 for flight.
+To enhance robustness, a variant of the datasets with added noise was considered, as described in the work from Panichi et al.
+(2025).
+Figure 4: Sequential images of quadruped jumps: (a) A simulated Go1 robot performs a pronking jump. (b) A simulated A1 executes a froggy jump. (c) A real Go1 performs a pronking jump.
+Real-world data were collected from a Go1 robot, as depicted in Fig.
 4(c).
@@
 3.4 Evaluation Procedure
+To validate the model $\Theta$, the initial state $(\mathbf{q}_{0},\dot{\mathbf{q}}_{0})$ and inputs $\mathbf{u}$ are transformed by the trained encoder and the input transformation equation to the latent initial state $(\bm{\xi}_{0},\dot{\bm{\xi}}_{0})$ and $\bm{\nu}$, respectively.
+The model then predicts latent accelerations $\ddot{\bm{\xi}}_{\text{pred}}$, which are integrated using the lsoda method (Hindmarsh and Petzold, 2005) to obtain predicted latent space state ${\bm{\xi}}_{\text{pred}}$ and $\dot{\bm{\xi}}_{\text{pred}}$ with a step size of $500\text{\,}\mathrm{H}\mathrm{z}$.
+This process is repeated for each time step.
+The predicted latent configuration $\bm{\xi}_{\text{pred}}$ is then decoded into a original configuration space by $\mathbf{q}_{\text{pred}}=\mathbf{W}_{\mathrm{d}}\,\bm{\xi}_{\text{pred}}+\mathbf{B}_{\mathrm{d}}$ for direct comparison with actual jump data.
+An alternative evaluation approach involves simulating the learned system for shorter intervals (e.g., $0.1\text{\,}\mathrm{s}$ – 50 steps) with periodic resets to the original latent state $(\bm{\xi},\dot{\bm{\xi}})$.
+This strategy decreases error accumulation over time, showcasing the precision of the model for medium-horizon forecasts, which could be used in planning and control frameworks such as MPC.
+4 Results
+This section evaluates the approach.
+We explored several latent-space dimensionalities to assess the trade-off between model compactness and accuracy.
+Two representative configurations are reported in Table 1.
+Table 1: Hyperparameters of the models. The number of candidate functions is computed as $\sum_{k=0}^{p}(2l+k-1)+2(2l)+l$ for latent dimension $l$ and polynomial order pp.
+Parameter Symbol / Value Go1 A1
+Latent dimension $l$ 2 4
+Polynomial order pp 1 1
+# Candidates $|\Theta|$ 15 29
+Sparsity threshold $\lambda$ 0.01
+Learning rate $\eta$ $1\times 10^{-3}$
+Batch size $B$ 100
+Epochs $N_{\text{epoch}}$ 501
+4.1 Go1 Simulation Results
+4.1.1 Learning Go1 pronking
+We initially explored a two-dimensional latent model for the pronking.
+Fig.
+5 visualizes the transposed weight matrix $\mathbf{W}_{\mathrm{e}}^{\mathrm{T}}$ of the trained encoder, indicating that the latent dimensions primarily correspond to the CoM positions in the $x$ and $z$ coordinates.
+Figure 5: Heatmap of the encoder weight matrix $\mathbf{W}_{\mathrm{e}}^{\mathrm{T}}$ (Go1 simulation, two latent dimensions).
+Brighter colors indicate larger weights; rows are 18 configuration dimensions and columns are the two latent dimensions.
+The learned equations of motion in the latent space for the contact and flight phases, respectively, are
+$\begin{cases}\ddot{\xi}_{1}\!\!=\!\!0.36\!-\!0.16\,\dot{\xi}_{1}\!-\!\!0.91\,\dot{\xi}_{2}\!-\!0.75\sin(\dot{\xi}_{1})\!+\!0.11\,\nu_{1}\!-\!0.05\,\nu_{2},\\ \ddot{\xi}_{2}\!\!=\!\!-0.16\!-\!0.23\,\dot{\xi}_{1}\!-\!0.75\,\dot{\xi}_{2}\!-\!0.96\,\sin(\dot{\xi}_{2})\!+\!0.14\,\nu_{2},\end{cases}$ (12)
+and
+$\begin{cases}\ddot{\xi}_{1}=0.29-1.22\,\dot{\xi}_{1}-1.00\sin(\dot{\xi}_{1})+51.05\,\nu_{1},\\ \ddot{\xi}_{2}=0.42\,\dot{\xi}_{1}+0.52\sin(\dot{\xi}_{1}).\end{cases}$ (13)
+The latent coordinates do not correspond to explicit physical variables, but the learned symbolic equations display structures that are consistent with key principles of quadruped jumping.
+The constant terms behave like gravity-related biases, the linear velocity terms act as effective damping, and the input-dependent components reflect how actuation influences the motion.
+The trigonometric functions of latent velocities serve as smooth nonlinear shaping terms that encode phase-dependent effects arising from posture changes, contact transitions, and limb coordination.
+These nonlinear components naturally emerge when compressing high-dimensional multibody dynamics into a low-dimensional latent representation.
+Figure 6: Evaluation of the trained model (with two latent dimensions) for simulated Go1 pronking.
+Considering the symmetry between right and left side, we only report the joints angle profiles in front left (fl) and rear right ($r$) leg.
+Table 2: Prediction performance of the trained model on simulated and real datasets. Values denote mean absolute error $\pm$ standard deviation for joint angles and CoM pose components.
+Dataset Pre. Joint angles CoM position CoM orientation
+fl hx fl hy fl hz rr hx rr hy rr hz x y z roll pitch yaw
 […]
-In this study, we regard the aSLIP model (see Appendix.A) as the baseline for predicting the dynamical behavior.
-4 Results
-This section evaluates the approach using simulated Unitree Go1 and A1 quadrupeds data and real-world data from a Unitree Go1.
-4.1 Go1 Simulation Results
-4.1.1 Learning in 2D Latent Space for Go1
-We initially explored a two-dimensional latent model for the synchronized jumping.
+[…]
+[…]
+[…]
+[…]
+[…]
+The prediction results for Go1 pronking are shown in Fig.
+6.
+Table 2 reports the mean errors and the standard deviation (std) along each direction, demonstrating the high prediction accuracy.
+To furthermore demonstrate the effectiveness, we regard the actuated SLIP (aSLIP) model (see Appendix.A) as the baseline.
 Fig.
-5 visualizes the transposed weight matrix $\mathbf{W}_{\mathrm{e}}^{\mathrm{T}}$ of the trained encoder, indicating that the (learned) latent dimensions primarily correspond to the robot’s body center positions in the $x$ and $z$ coordinates.
-Figure 5: Heatmap of the encoder’s transposed weight matrix $\mathbf{W}_{\mathrm{e}}^{\mathrm{T}}$ trained on the simulated Go1 dataset with two latent dimensions.
-Brighter colors indicate higher absolute weights.
-The vertical axis shows the 18 configuration dimensions, and the horizontal axis shows the two latent dimensions.
-The learned equations of motion in the latent space for the contact and flight phases, respectively, are
-$\begin{cases}\ddot{\xi}_{1}\!\!=\!\!0.36\!-\!0.16\,\dot{\xi}_{1}\!-\!\!0.91\,\dot{\xi}_{2}\!-\!0.75\sin(\dot{\xi}_{1})\!+\!0.11\,\nu_{1}\!-\!0.05\,\nu_{2}\\ \ddot{\xi}_{2}\!\!=\!\!-0.16\!-\!0.23\,\dot{\xi}_{1}\!-\!0.75\,\dot{\xi}_{2}\!-\!0.96\,\sin(\dot{\xi}_{2})\!+\!0.14\,\nu_{2}\end{cases},$ (12)
-and
-$\begin{cases}\ddot{\xi}_{1}=0.29-1.22\,\dot{\xi}_{1}-1.00\sin(\dot{\xi}_{1})+51.05\,\nu_{1}\\ \ddot{\xi}_{2}=0.42\,\dot{\xi}_{1}+0.52\sin(\dot{\xi}_{1})\end{cases}.$ (13)
-The prediction results for simulated Go 1 jumping are shown in Fig. 6.
-Figure 6: Evaluation of the trained model with two latent dimensions on the simulated Go1 jumping dataset.
-Fig. 7 compares the body’s linear position along the $x$ and $z$ axes, demonstrating that the learned model outperforms, especially in the $x$ axis, the baseline aSLIP model.
-Figure 7: Comparison between the learned model and the aSLIP model, on the task of predicting the CoM linear positions along the $x$ and $z$ axes, where the simulated Go 1 performs a jumping motion.
+7 compares the body’s linear position along the $x$ and $z$ axes, demonstrating that the learned model outperforms the baseline reduced-order model.
+Statistical results demonstrate that the average prediction error is reduced from 5.7e-2 m to 1.6e-2 m along the $x$ direction and reduced from 9.5e-2 m to 4.0e-2 m along $z$ direction.
+Figure 7: Comparison between the learned model and the aSLIP model regarding the CoM positions prediction, where the Go 1 performs a pronking jumping motion.
 4.1.2 Latent Space Dimensionality Tradeoff Analysis
-We analyze the effect of latent dimensionality on model design. As shown in Fig. 8, increasing to 18 dimensions does not change the dominance of key full-order components.
-Figure 8: Heatmap visualization of the transposed weight matrix $\mathbf{W}_{\mathrm{e}}^{\mathrm{T}}$ of the encoder trained on the simulated Go1 dataset in the case of 181818 latent dimensions.
-Fig.9 further evaluates the trade-off between accuracy (via final reconstruction error $\mathcal{E}_{\text{dec}}$) and complexity (via active SINDy coefficients $|\bm{\Xi}|$) across 18 contact phase models with varying latent dimensions.
-Figure 9: The trade-off between accuracy and complexity with increasing latent dimensions evaluated on the simulated Go1 dataset.
-The error bars depict the standard deviation from five different random seeds.
-To quantitatively evaluate this trade-off, we employ a combined loss $\mathcal{L}_{\text{mod}}$ that is inspired by the Akaike Information Criterion (AIC) for model selection (Bozdogan, 1987) and calculated as
+Fig. 8 illustrates the effect of latent dimensionality on model design.
+Figure 8: Heatmap of the encoder weight matrix $\mathbf{W}_{\mathrm{e}}^{\mathrm{T}}$ (Go1 simulation, 18 latent dimensions).
+Fig. 9 further evaluates the trade-off between accuracy (via final reconstruction error $\mathcal{E}_{\text{dec}}$) and complexity (via active coefficients $|\bm{\Xi}|$) across 18 latent dimensions.
+Figure 9: Accuracy vs. complexity with the increase of latent space dimensions.
+To quantitatively evaluate this trade-off, we employ a combined loss $\mathcal{L}_{\text{mod}}$ (Bozdogan, 1987), calculated as
 $\mathcal{L}_{\text{mod}}=2\,\mathcal{E}_{\text{dec}}+2\,\lambda\,\log{(|\bm{\Xi}|)},$ (14)
-with $\lambda=0.001$, balancing the loss terms, the combined loss (red line in Fig.9) suggests optimal performance with 4–6 latent dimensions.
-This aligns with the visibly significant components in Fig.8, where further increasing dimensionality yields marginal gains at the cost of complexity.
+with $\lambda=0.001$.
+The combined loss (red line in Fig.
+9) suggests optimal performance with 4–6 latent dimensions.
+This aligns with the visibly significant components in Fig.
+8, where further increasing dimensionality yields marginal gains at the cost of complexity.
 4.2 A1 Simulation Results
-We also conducted experiments on data collected with a simulated Unitree A1 quadruped executing a froggy jump to evaluate the algorithm’s effectiveness across different platforms.
-The results are illustrated in Fig.
-10.
-Figure 10: Evaluation of the proposed method with four latent dimensions on A1 froggy jumping.
-A notable difference is observed in the body angular position $y$ between the Unitree A1 and the Unitree Go1.
+We also conducted experiments with a simulated A1 froggy jump to evaluate the algorithm’s effectiveness across different platforms and modalities. The results are illustrated in Fig. 10 and Table 2.
+Figure 10: Learning A1 froggy jumping (simulation data) with four latent dimensions.
+A notable difference is observed in the body angular position $y$ (i.e., pitch angle) between the Unitree A1 and Go1.
 The froggy jump causes a more pronounced angular rotation along the $y$-axis in the A1’s case.
-Despite the increased significance of this angular dimension for the jump motion, the four-dimensional model successfully captures this dynamic and most joint positions.
+Despite the increased significance of this angular dimension for the froggy jump, the four-dimensional model successfully captures this dynamic and most joint positions.
+High precision is obtained when reset the latent state every 0.1 s.
 4.3 Go1 Real World Results
@@
 Fig.
-11 illustrates the model’s ability to replicate a jump on the hardware, comparing joint angles and CoM positions and orientations.
-Figure 11: Evaluation of the learned reduced-order model with four latent dimensions on the Go1 real-world testing dataset.
-The proposed model performs well on actual hardware data, despite reduced accuracy during flight and initial landing phases compared to simulations.
-Nonetheless, the model effectively manages real-world noise and complexity, demonstrating robust performance.
+11 illustrates the model’s ability to replicate a jump on the hardware, and Table 2 reports the prediction error.
+Figure 11: Learning Go1 pronking jumping (real data) with four latent dimensions.
+As can be seen from Fig.
+11 and Table 2, the proposed model performs well on actual hardware data, despite reduced accuracy during flight and initial landing phases compared to simulations.
+The divergence between real and predicted trajectories mainly results from unmodeled actuator and contact effects that accumulate during real execution.
+Applying a domain randomization in the future work strategy further alleviates this discrepancy by enhancing the model’s robustness to such unmodeled variations.
+The symbolic model still captures the key motion patterns, with residual errors from friction, delay, and noise remaining bounded.
 5 Discussion and Conclusion
-The results of our symbolic learning framework show that a reduced latent space can effectively preserve key dynamic patterns of the quadruped robot jumping motion.
-By combining autoencoder-based embeddings with sparse regression, we obtain interpretable and compact models that approximate the complex full-order dynamics.
-Our analysis of latent dimensionality shows that increasing beyond 4–6 dimensions offers diminishing returns in accuracy while introducing additional model complexity.
-The sim-to-real transfer results indicate that the model retains core dynamic features and generalizes well despite limited real-world data, validating its robustness to measurement noise and domain shift.
-In future work, we aim to extend this methodology to more dynamic behaviors, such as walking and running, and to systems involving rapid contact changes or higher-dimensional morphologies, including humanoids.
+This work demonstrates that a low-dimensional, physics-aligned latent space can faithfully capture the essential dynamics of quadruped jumping.
+By combining a linear autoencoder embedding with sparse regression, the proposed framework yields compact and interpretable symbolic models that approximate full-order dynamics with meaningful accuracy.
+The learned equations recover gravity-like loading, effective damping, and actuation influence, while revealing coupling effects absent in idealized template models yet necessary to represent the richer behaviors of the full quadruped system.
+Beyond these empirical results, this study highlights how a linear latent embedding—together with a coherent multi-phase training strategy—enables SINDy to discover reduced-order template dynamics for hybrid legged motion, marking a conceptual departure from prior nonlinear SINDy-AE frameworks.
+The approach remains sensitive to unmodeled actuator and contact effects, which accumulate during long-horizon rollouts and partly explain the observed sim-to-real discrepancies.
+Future work will extend the methodology to more diverse locomotion behaviors, incorporate hybrid contact transitions more explicitly, and scale the framework to higher-dimensional morphologies.
+Another promising direction is to integrate the learned symbolic models into adaptive or phase-dependent controllers for real-time robot movement tasks.
 Appendix A Baseline: actuated SLIP Model